Electronic Journal: Southwest Journal of Pure and Applied Mathematics Internet: http://rattler.cameron.edu/swjpam.html ISSN 1083-0464 Vol. No. 02, 1996, pp. 60–64 Submitted: October 17, 1996. Published: December 9, 1996.

A GENERALIZATION OF EDELSTEIN’S THEOREM

ON FIXED POINTS AND APPLICATIONS

Ioannis K. Argyros

Abstract. In this study, we approximate fixed points of nonlinear equations on topological vector spaces using implicit methods. A new Edelstein type theorem on fixed points is proved. The advantage of these methods is that they can be used to solve equations that usual explicit methods cannot.

(1991) A.M.S. Subject Classification Codes. 65J15, 47H17, 49D15.

Key Words and Phrases. Implicit methods, fixed point, topological vector space, normed space.

Typeset by AMS-TEX

Cameron University, Lawton, Oklahoma

E-mail Address: ioannisa@cameron.edu

\begin{matrix}c \\ ◯ \end{matrix}1996 Cameron University